# compat=1.0.0

### v1

@@ 1 Introduction
To address the challenges related to the origin of neutrino masses, their flavor structure, the presence of two distinct mass scales, and the stability of DM, a recent approach based on flavor symmetries has been developed [26, 27, 28, 29]. Within this framework, it has been demonstrated that employing the $A_{4}$ flavor symmetry [30, 31, 32] allows for predictive solutions to the leptonic flavor structure. In addition, the spontaneous breaking of $A_{4}$ to its subgroup $\mathcal{Z}_{2}$ [33, 34, 35] ensures DM stability while also leading to the scoto-seesaw mass mechanism. Refs. [26, 27], explore the Majorana scotogenic loop, whereas Ref. [28] investigates the Dirac scotogenic loop. Additionally, the authors of [27] emphasize that the mass of DM is tightly constrained by the breaking of $A_{4}$ symmetry. They further provide conclusive evidence for the scalar DM scenario by ruling out the possibility of fermionic DM.
@@
Here, we also consider an accidental $U(1)_{B-L}$ symmetry, inherently present in the SM, and its breaking into a $\mathcal{Z}_{3}$ subgroup to prevent tree-level couplings before SSB and to forbid the Majorana neutrino mass term. The chiral anomaly-free $B-L$ symmetry requires the ($-4,-4,5$) charge assignment for the right-handed neutrinos $\nu_{R_{i}}$, which makes the lightest neutrino massless. Unlike the Majorana version [26, 27], our model features both normal ordering (NO) and inverted ordering (IO) of neutrino masses. Furthermore, the $A_{4}$ symmetric approach imposes an upper bound on the scalar masses based on the two mass-squared differences of neutrino, thereby constraining the allowed parameter space for DM candidates. Also, we study the different charged lepton flavor violation (CLFV) processes and check their viability under different current and future proposed experiments.
@@ 2 Model SET-UP
The charge assignments in Tab. 1 ensure that, following the $A_{4}$ symmetry breaking by the VEVs of $\eta$ and $\xi$, all the SM particles and $\nu_{R_{i}}$’s remain even under the residual $\mathcal{Z}_{2}$ symmetry. In contrast, the triplet BSM particles may acquire odd charges under the residual $\mathcal{Z}_{2}$ symmetry. The components of $A_{4}$ triplets $\eta,\ \xi$ and $N$ split into two, the first components transforming as even while the other components have odd charges under $\mathcal{Z}_{2}$, see App. B for more details. The odd particles will eventually become dark sector particles as discussed in Sec. 1. This discussion will focus on the scenario, where $U(1)_{B-L}$ breaks down to the residual $\mathcal{Z}_{3}$ subgroup [^1] The explicit breaking of the $U(1)_{B-L}$ symmetry arises from the cubic interaction term $H^{\dagger}\eta\,\xi$ in the scalar potential. Here, $H$ and $\xi$ carry zero $B-L$ charge, while $\eta$ has a charge of three. Consequently, this interaction alters $B-L$ in steps of three, leading to the breaking of $U(1)_{B-L}$ in units of three. As a result, the residual symmetry is the discrete subgroup $\mathcal{Z}_{3}$. Since $B-L$ is violated in multiples of three, the transformations that preserve the term $H^{\dagger}\eta\,\xi$ must correspond to those where $B-L$ charges differ by multiples of three. This implies that the $U(1)_{B-L}$ phase factor, $e^{i\alpha(B-L)}$, is constrained such that $e^{i\alpha(B-L)}=e^{i\alpha(B-L+3n)}$ for any integer $n$. This periodicity signifies that the continuous $U(1)_{B-L}$ symmetry is broken down to the discrete subgroup $\mathcal{Z}_{3}$., a natural outcome when the $B-L$ symmetry is broken in a unit of three. The presence of the residual $\mathcal{Z}_{3}$ subgroup ensures the Dirac nature of neutrinos.
@@
  • In this minimal setup, one neutrino remains massless because the Yukawa interaction $Y^{\prime}\bar{N}_{L}\xi\nu_{R}$ is forbidden for the third-generation right-handed neutrino due to $B-L$ charge assignments. A Dirac mass can be introduced by adding an $SU(2)_{L}$ singlet but $A_{4}$ triplet scalar $\zeta_{9}$ with $B-L$ charge $-9$, allowing the Yukawa term $y_{9}\bar{N}_{L}\zeta_{9}\nu_{R_{3}}$. After SSB, if $\zeta_{9}$ acquires a VEV $(u_{9},0,0)$, it generates a Dirac mass for $\nu_{R_{3}}$. Since $\zeta_{9}$ is an $SU(2)_{L}$ singlet, its VEV can be much larger than the electroweak scale, giving $\nu_{R_{3}}$ a large mass and effectively decoupling both $\nu_{R_{3}}$ and $\zeta_{9}$ from the rest of the particles.
@@ 2.3 Neutrino Masses
At the tree level, the $(6\times 6)$ mass matrix in the basis $(\overline{\nu}_{L_{i}},\overline{N}_{L_{j}})$ and $(\nu_{R_{i}},N_{R_{j}})^{T}$, (where $i\,,j=1,2,3$) is given by
@@
$-m^{(1)}_{\nu}=\mathcal{M}_{D}\mathcal{M}^{-1}\boldsymbol{\mu}=\frac{v_{2}u}{\sqrt{2}M}\begin{pmatrix}y_{1}y^{\prime}_{1}&y_{1}y^{\prime}_{2}&0\\ y_{2}y^{\prime}_{1}&y_{2}y^{\prime}_{2}&0\\ y_{3}y^{\prime}_{1}&y_{3}y^{\prime}_{2}&0\\ \end{pmatrix}\;.$ (17)
The resulting neutrino mass is proportional to the $\boldsymbol{\mu}$-matrix, where $u$ is the VEV of the singlet field $\xi$. Since $u$ can take very small values, this feature is characteristic of the inverse seesaw mechanism [37], which will be discussed in the next section. It is noteworthy that the rank of $m^{(1)}_{\nu}$ is one, implying that only one of the neutrinos acquires mass at the tree level. The additional neutrino masses can be generated at the loop level from the scoto-loop mechanism. The loop contribution is given by
@@ 3.2 Inverted Ordering
For IO, it is noteworthy that, unlike the Majorana case [26, 27], where IO was forbidden, in this Dirac neutrino formalism, IO is permitted and remains consistent with the observed neutrino oscillation data. As in IO, neutrino mass ordering follows $m_{2}>m_{1}>m_{3}$, hence the lightest one, $m_{3}=0$ for this case. Imposing $\Delta m^{2}_{ij}$ constraint, we find $m_{1}$ and $m_{2}$ are as follows:
@@ 4 Dark Sector Results
$\lambda_{1}$ $[10^{-4},\sqrt{4\pi}]$ $\lambda_{\eta_{l}}$ $[10^{-8},\sqrt{4\pi}]$
$\lambda_{\xi_{k}}$ $[10^{-8},\sqrt{4\pi}]$ $|\lambda_{H\eta_{i}}|$ $[10^{-8},\sqrt{4\pi}]$
$|\lambda_{H\xi_{m}}|$ $[10^{-8},\sqrt{4\pi}]$ $|\lambda_{\eta\xi_{n}}|$ $[10^{-8},\sqrt{4\pi}]$
$|\kappa|$ $[10^{-8},30]\text{ GeV}$ $u$ $[10^{-4},10^{5}]\text{ GeV}$
$M_{N_{i}}$ $[10,10^{12}]\text{ GeV}$ $y_{i}$, $y^{\prime}_{j}$ $[10^{-6},10^{-1}]$
@@ 5 Charged Lepton Flavor violation
The discovery of neutrino oscillations provides compelling evidence for the existence of charged lepton flavor violating interactions [49]. In the SM, if neutrino masses originate similarly to those of other fermions (via Yukawa interactions with the SM Higgs), charged lepton flavor violating (CLFV) rates are expected to be extremely suppressed to practically unobservable levels. For instance, the branching fraction for $\mu\rightarrow e\gamma$ decay is estimated to be of the order of $10^{-54}$, significantly below the sensitivity threshold of any feasible experiment. However, alternative mechanisms for neutrino mass generation could yield to potentially significant CLFV effects. Extensive experimental efforts have been made to explore flavor-violating decays of muons and tau leptons. Yet, as of the present date, no conclusive positive signals have emerged from these experiments. Consequently, the absence of such signals has prompted the establishment of stringent bounds on the branching ratios associated with these flavor-violating decay processes. Presently, these limits stand at $BR(\mu\rightarrow e\gamma)<3.1\times 10^{-13}$ [50], $BR(\tau\rightarrow e\gamma)<3.3\times 10^{-8}$ [51], and $BR(\tau\rightarrow\mu\gamma)<4.2\times 10^{-8}$ [52], respectively. Projections indicate significant enhancements in the near future for detecting these decay processes, with anticipated limits of $BR(\mu\rightarrow e\gamma)<6\times 10^{-14}$ [53], $BR(\tau\rightarrow e\gamma)<9\times 10^{-9}$ [54], and $BR(\tau\rightarrow\mu\gamma)<6.9\times 10^{-9}$ [54], respectively. The leading contribution to $\ell_{\alpha}\rightarrow\ell_{\beta}\gamma$ arises at the one-loop level through the mediation of the charged scalar $\eta^{+}_{i}$ as shown in the diagram of Fig. 11. The branching ratio of this process is given by
@@
We have analyzed various CLFV rates corresponding to different DM candidates. In Fig. 12, we show the resulting CLFV decay rates as a function of the mass for each viable DM candidate in our model. The gray and olive green points represent regions of under-abundance and over-abundance, respectively, while the blue points correspond to the correct relic density. Note that all three processes show similar behavior: their maximum values drop as the DM mass increases. This behavior arises because DM is the lightest particle in the dark sector. As its mass increases, the other dark sector particles also become heavier, which lowers the CLFV rates, as can be seen from Eq. (31).
@@
In our analysis, we identified viable data points in the mass ranges of (63–150) GeV for the doublet scalar, (48–265) GeV for the singlet scalar, and (50–160) GeV for the fermionic DM candidate, all of which satisfy the previously discussed experimental constraints. Notably, these points lie between the current experimental bounds and the projected future sensitivity for the decay $\mu\rightarrow e\gamma$. However, for the decay channels $\tau\rightarrow e\gamma$ and $\tau\rightarrow\mu\gamma$, no points within our scanned parameter space fall within the current or expected experimental limits. This is primarily due to the comparatively weaker experimental bounds for these decay modes, which our model does not reach within the explored parameter range. Thus, CLFV points that are both within experimental reach and consistent with DM constraints appear only in a few distinct regions, making CLFV measurements a valuable probe for testing our model.
@@ 6 Conclusions
We propose a Dirac scotogenic-like model where DM stability arises from the breaking of an $A_{4}$ flavor symmetry. The same breaking also induces a “scoto inverse-seesaw” mechanism, which combines a tree-level inverse-seesaw with a one-loop scotogenic contribution to generate neutrino masses. The inverse seesaw mechanism at the tree level is enabled by the smallness of the singlet VEV $u$. This framework naturally explains the observed neutrino mass-squared differences, $\Delta m^{2}_{\rm{atm}}$ and $\Delta m^{2}_{\rm{sol}}$, from oscillation data. The model also addresses leptonic flavor patterns via the $A_{4}$ symmetry. It includes right-handed neutrinos $\nu_{R}$, $SU(2)_{L}$ singlet fermions $N_{L}$, $N_{R}$, a singlet scalar $\xi$, and a doublet scalar $\eta$. Here, $\nu_{R}$’s are $A_{4}$ singlets, while all other BSM particles transform as $A_{4}$ triplets. To forbid tree-level neutrino couplings before symmetry breaking, we assign specific $B-L$ charges $(-4,-4,5)$ to $\nu_{R}$. The breaking $U(1)_{B-L}\to\mathcal{Z}_{3}$ then ensures neutrinos remain Dirac in nature and results in one massless neutrino.

### v2

# Dirac Scoto Inverse-Seesaw from $A_{4}$ Flavor Symmetry

@@ 1 Introduction
To address the challenges related to the origin of neutrino masses, their flavor structure, the presence of two distinct mass scales, and the stability of DM, a recent approach based on flavor symmetries has been developed [26, 27, 28]. Within this framework, it has been demonstrated that employing the $A_{4}$ flavor symmetry [29, 30, 31] allows for predictive solutions to the leptonic flavor structure. In addition, the spontaneous breaking of $A_{4}$ to its subgroup $\mathcal{Z}_{2}$ [32, 33, 34] ensures DM stability while also leading to the scoto-seesaw mass mechanism. Refs. [26], explore the Majorana scotogenic loop, whereas Ref. [27] investigates the Dirac scotogenic loop. Additionally, the authors of [26] emphasize that the mass of DM is tightly constrained by the breaking of $A_{4}$ symmetry. They further provide conclusive evidence for the scalar DM scenario by ruling out the possibility of fermionic DM.
@@
Here, we also consider an accidental $U(1)_{B-L}$ symmetry, inherently present in the SM, and its breaking into a $\mathcal{Z}_{3}$ subgroup to prevent tree-level couplings before SSB and to forbid the Majorana neutrino mass term. The chiral anomaly-free BLB-L symmetry requires the (4,4,5-4,-4,5) charge assignment for the right-handed neutrinos $\nu_{R_{i}}$ [37, 38, 39]. Unlike the Majorana version [26], our model features both normal ordering (NO) and inverted ordering (IO) of neutrino masses. Furthermore, the $A_{4}$ symmetric approach imposes an upper bound on the scalar masses based on the two mass-squared differences of neutrino, thereby constraining the allowed parameter space for DM candidates. Also, we study the different charged lepton flavor violation (CLFV) processes and check their viability under different current and future proposed experiments.
@@ 2 Model SET-UP
The charge assignments in Tab. 1 ensure that, following the $A_{4}$ symmetry breaking by the VEVs of $\eta$ and $\xi$, all the SM particles and $\nu_{R_{i}}$’s remain even under the residual $\mathcal{Z}_{2}$ symmetry. In contrast, the triplet BSM particles may acquire odd charges under the residual $\mathcal{Z}_{2}$ symmetry. The components of $A_{4}$ triplets $\eta,\ \xi$ and $N$ split into two, the first components transforming as even while the other components have odd charges under $\mathcal{Z}_{2}$, see App. B for more details. The odd particles will eventually become dark sector particles as discussed in Sec. 1. This discussion will focus on the scenario, where $U(1)_{B-L}$ breaks down to the residual $\mathcal{Z}_{3}$ subgroup [^1] The explicit breaking of the $U(1)_{B-L}$ symmetry arises from the cubic interaction term $H^{\dagger}\eta\,\xi$ in the scalar potential. Here, $H$ and $\xi$ carry zero BLB-L charge, while $\eta$ has a charge of three. Consequently, this interaction alters BLB-L in steps of three, leading to the breaking of $U(1)_{B-L}$ in units of three. As a result, the residual symmetry is the discrete subgroup $\mathcal{Z}_{3}$. Since BLB-L is violated in multiples of three, the transformations that preserve the term $H^{\dagger}\eta\,\xi$ must correspond to those where BLB-L charges differ by multiples of three. This implies that the $U(1)_{B-L}$ phase factor, $e^{i\alpha(B-L)}$, is constrained such that $e^{i\alpha(B-L)}=e^{i\alpha(B-L+3n)}$ for any integer $n$. This periodicity signifies that the continuous $U(1)_{B-L}$ symmetry is broken down to the discrete subgroup $\mathcal{Z}_{3}$., a natural outcome when the BLB-L symmetry is broken in a unit of three. The presence of the residual $\mathcal{Z}_{3}$ subgroup ensures the Dirac nature of neutrinos.
@@
  • In this minimal setup, one neutrino remains massless because the Yukawa interaction $Y^{\prime}\bar{N}_{L}\xi\nu_{R}$ is forbidden for the third generation right-handed neutrino due to BLB-L charge assignments. However, the massless neutrino can be made massive by adding a $SU(2)_{L}$ singlet but $A_{4}$ triplet scalar $\zeta_{9}$ with BLB-L charge 9-9, allowing the Yukawa term $y_{9}\bar{N}_{L}\zeta_{9}\nu_{R_{3}}$. After SSB, $\zeta_{9}$ acquires a VEV $(u_{9},0,0)$ preserving the $A_{4}\to\mathcal{Z}_{2}$ breaking. Then through the scoto-seesaw mechanism, we can generate a light Dirac neutrino mass for the $(\nu_{L_{3}},\nu_{R_{3}})$ pair. Note that, since one massless neutrino is allowed by the current oscillation data, we do not develop this analysis further.
@@ 2.3 Neutrino Masses
At the tree level, the $(6\times 6)$ mass matrix in the basis $(\overline{\nu}_{L_{i}},\overline{N}_{L_{j}})$ and $(\nu_{R_{i}},N_{R_{j}})^{T}$, (where i,j=1,2,3i\,,j=1,2,3) is given by
@@
$m^{(1)}_{\nu}=\mathcal{M}_{D}\mathcal{M}^{-1}\boldsymbol{\mu}=\frac{v_{2}u}{\sqrt{2}M}\begin{pmatrix}y_{1}y^{\prime}_{1}&y_{1}y^{\prime}_{2}&0\\ y_{2}y^{\prime}_{1}&y_{2}y^{\prime}_{2}&0\\ y_{3}y^{\prime}_{1}&y_{3}y^{\prime}_{2}&0\\ \end{pmatrix}\;$. (17)
The resulting neutrino mass is proportional to the $\boldsymbol{\mu}$-matrix, where $u$ is the VEV of the singlet field $\xi$. Since $u$ can take very small values, this feature is characteristic of the inverse seesaw mechanism [36], which will be discussed in the next section. It is noteworthy that the rank of $m^{(1)}_{\nu}$ is one, implying that only one of the neutrinos acquires mass at the tree level.
Note that, even when we take the one-loop contribution associated with $N_{1}$ into account, the texture zeros of Eq. (17) remain symmetry protected. This results in an additional mass matrix with a similar structure to Eq. (17). It is due to the fact that the residual $\mathcal{Z}_{2}$ symmetry remains unbroken, and as a consequence, mixing between the first generation singlet fermion $N_{1}$ (even under $\mathcal{Z}_{2}$) and the other generations $N_{2}$, $N_{3}$ (odd under $\mathcal{Z}_{2}$) is forbidden. Similarly, mixing between scalars $\eta_{1}$ or $\xi_{1}$ (even under $\mathcal{Z}_{2}$) and the dark scalars $\eta_{2,3}$ or $\xi_{2,3}$ (odd under $\mathcal{Z}_{2}$) are also not allowed by this residual symmetry. As a result, the loop correction do not induce any new non-zero entries in the neutrino mass matrix; the zero entries remain protected by symmetry. The loop induced corrections affect only the non-zero elements of the tree-level mass matrix in Eq. (17), which can be effectively absorbed into a redefinition of the Yukawa couplings. We do not explicitly include these loop contribution in our analysis. Thus, the overall structure of the mass matrix remains unchanged and only one neutrino mass can be generated at this stage.
The additional neutrino masses can be generated at the loop level from the scoto-loop mechanism. The loop contribution is given by
@@ 3.2 Inverted Ordering
For IO, it is noteworthy that, unlike the Majorana case [26], where IO was forbidden, in this Dirac neutrino formalism, IO is permitted and remains consistent with the observed neutrino oscillation data. As in IO, neutrino mass ordering follows $m_{2}>m_{1}>m_{3}$, hence the lightest one, $m_{3}=0$ for this case. Imposing $\Delta m^{2}_{ij}$ constraint, we find $m_{1}$ and $m_{2}$ are as follows:
@@ 4 Dark Sector Results
$\lambda_{1}$ $10^{-4},\sqrt{4\pi}]$ $\lambda_{\eta_{l}}$ $10^{-8},\sqrt{4\pi}]$
$\lambda_{\xi_{k}}$ $10^{-8},\sqrt{4\pi}]$ $|\lambda_{H\eta_{i}}|$ $10^{-8},\sqrt{4\pi}]$
$|\lambda_{H\xi_{m}}|$ $10^{-8},\sqrt{4\pi}]$ $|\lambda_{\eta\xi_{n}}|$ $10^{-8},\sqrt{4\pi}]$
$|\kappa|$ $10^{-8},30]\text{ GeV}$ $u$ $10^{-4},10^{5}]\text{ GeV}$
$M_{N_{i}}$ $10,10^{12}]\text{ GeV}$ $y_{i}$, $y^{\prime}_{j}$ $10^{-6},10^{-1}]$
@@ 5 Charged Lepton Flavor violation
The discovery of neutrino oscillations provides compelling evidence for the existence of charged lepton flavor violating interactions [51]. In the SM, if neutrino masses originate similarly to those of other fermions (via Yukawa interactions with the SM Higgs), charged lepton flavor violating (CLFV) rates are expected to be extremely suppressed to practically unobservable levels. For instance, the branching fraction for $\mu\rightarrow e\gamma$ decay is estimated to be of the order of $10^{-54}$, significantly below the sensitivity threshold of any feasible experiment. However, alternative mechanisms for neutrino mass generation could yield to potentially significant CLFV effects. Extensive experimental efforts have been made to explore flavor-violating decays of muons and tau leptons. Yet, as of the present date, no conclusive positive signals have emerged from these experiments. Consequently, the absence of such signals has prompted the establishment of stringent bounds on the branching ratios associated with these flavor-violating decay processes. Presently, these limits stand at $BR(\mu\rightarrow e\gamma)<3.1\times 10^{-13}$ [52], $BR(\tau\rightarrow e\gamma)<3.3\times 10^{-8}$ [53], and $BR(\tau\rightarrow\mu\gamma)<4.2\times 10^{-8}$ [54], respectively. Projections indicate significant enhancements in the near future for detecting these decay processes, with anticipated limits of $BR(\mu\rightarrow e\gamma)<6\times 10^{-14}$ [55], $BR(\tau\rightarrow e\gamma)<9\times 10^{-9}$ [56], and $BR(\tau\rightarrow\mu\gamma)<6.9\times 10^{-9}$ [56], respectively. The leading contribution to $\ell_{\alpha}\rightarrow\ell_{\beta}\gamma$ arises at the one-loop level through the mediation of the charged scalar $\eta^{+}_{i}$ as shown in the diagram of Fig. 11.
The branching ratio of this process is given by
@@
We have analyzed various CLFV rates corresponding to different DM candidates. In Fig. 12, we show the resulting CLFV decay rates as a function of the mass for each viable DM candidate in our model.
@@
The gray and olive green points represent regions of under-abundance and over-abundance, respectively, while the blue points correspond to the correct relic density. Note that all three processes show similar behavior: their maximum values drop as the DM mass increases. This behavior arises because DM is the lightest particle in the dark sector. As its mass increases, the other dark sector particles also become heavier, which lowers the CLFV rates, as can be seen from Eq. (31). In our analysis, we identified viable data points in the mass ranges of (63–150) GeV for the doublet scalar, (48–265) GeV for the singlet scalar, and (50–160) GeV for the fermionic DM candidate, all of which satisfy the previously discussed experimental constraints. Notably, these points lie between the current experimental bounds and the projected future sensitivity for the decay $\mu\rightarrow e\gamma$. However, for the decay channels $\tau\rightarrow e\gamma$ and $\tau\rightarrow\mu\gamma$, no points within our scanned parameter space fall within the current or expected experimental limits. This is primarily due to the comparatively weaker experimental bounds for these decay modes, which our model does not reach within the explored parameter range. Thus, CLFV points that are both within experimental reach and consistent with DM constraints appear only in a few distinct regions, making CLFV measurements a valuable probe for testing our model.
@@ 6 Conclusions
We propose a Dirac scotogenic-like model where DM stability arises from the breaking of an $A_{4}$ flavor symmetry. The same breaking also induces a “scoto inverse-seesaw” mechanism, which combines a tree-level inverse-seesaw with a one-loop scotogenic contribution to generate neutrino masses. The inverse seesaw mechanism at the tree level is enabled by the smallness of the singlet VEV $u$. This framework naturally explains the observed neutrino mass-squared differences, $\Delta m^{2}_{\rm{atm}}$ and $\Delta m^{2}_{\rm{sol}}$, from oscillation data. The model also addresses leptonic flavor patterns via the $A_{4}$ symmetry. It includes right-handed neutrinos $\nu_{R}$, $SU(2)_{L}$ singlet fermions $N_{L}$, $N_{R}$, a singlet scalar $\xi$, and a doublet scalar $\eta$. Here, $\nu_{R}$’s are $A_{4}$ singlets, while all other BSM particles transform as $A_{4}$ triplets. To forbid tree-level neutrino couplings before symmetry breaking, we assign specific BLB-L charges (4,4,5)(-4,-4,5) to $\nu_{R}$. The breaking $U(1)_{B-L}\to\mathcal{Z}_{3}$ then ensures neutrinos remain Dirac in nature and results in one massless neutrino.